AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 10 of 33  ELOG logo
ID Date Author Subject
  485   Thu Feb 1 14:45:09 2024 JM, CBNewMerger setup

To include alll 16 FEE modules in the NewMerger we edited the file /MIDAS/DB/EXPERIMENTS/MERGE/Options/carme-gsi from MERGE.LinksAvailable string 5 to MERGE.LinksAvailable string 15

We also edited the file /MIDAS/Linux/startup/NewMerger from ./master64 -i 8 -l 16 -p 11001  & to ./master64 -i 16 -l 16 -p 11001  &

We have started the Merger and tape server. The merger settings were 'corrupted' and did not display xfer Links => Merger. We implemented a fix from https://elog.ph.ed.ac.uk/DESPEC/36 and this resolved the issue.

We started the tape server, merger and go on the run control. We observe the merger working and data links to the tape server. A new file is created each time but we observe no data being written to that file. After 5 minutes of no data being written the merger crashed and the fees crashed soon after. We will leave until tomorrow to resolve.

Issue subsequently resolved see elog https://elog.ph.ed.ac.uk/CARME/487

  484   Thu Feb 1 10:04:11 2024 JM, CBShaping time testing

Arrived at the control room to find we had lost connection to the power relay. We rebooted the nnrpi1 and re-connected according to the steps laid out in elog 471 -> all okay

11:00 Restarted daq with no changes to the system from yesterday. Bias consistent with previous days. All pulser peaks are slightly improved (even more for the pn pulser, one goes from 91 chn to 78). No pulser observed in nn FEE modules 9, 10, 13, 14 again.

ADC data items check OK - all ~3k

Rates - all OK  - attachment 1

1.8L spectra. - attachment 2,3

1.8W waveforms - attachment 4,5

11:20 Changed shaping time from 0xf to 0x3 (2us)

ADC data items check OK - all ~3k attachment 6

Rates - all OK  - attachment 7

1.8L pn spectra. - attachment 8

1.8W waveforms - attachment 9,10

11:35 Changed shaping time from 0x3 to 0x7 (4us)

ADC data items check OK - all ~3k attachment 11

Rates - all OK  - attachment 12

1.8L pn spectra. - attachment 13

1.8W waveforms - attachment 14,15

11:45 Changed shaping time from 0x7 to 0xb (6us)

ADC data items check OK - all ~3k - attachment 16

Rates - all OK  - attachment 17

1.8L spectra. - attachment 18,19

1.8W waveforms - attachment 20,21

11:59 Changed shaping time from 0xb to 0xd (7us)

ADC data items check OK - all ~3k - attchment 22

Rates - all OK  - attachment 23

1.8L pn spectra. - attachment 24

1.8W waveforms - attachment 25, 26

12:09 Changed shaping time back to 0xf (8us)

ADC data items check OK - all ~3k - attchment 27

Rates - all OK  - attachment 28

1.8L pn spectra. - attachment 29

1.8W waveforms - attachment 30,31

13:10 Power cycled FEE modules. nn pulser for FEEs 9,10,13,14 are now active again (no changes made in the cave) Pulser widths for nn and pn  both consistent with before lunch

 

  483   Wed Jan 31 09:49:38 2024 JM, TDOptimising - Wednesday 31

Removed ground from DSSD #1 to the chamber and set MACBs so that timestamp is from the NIM WR emulator not the VME VETAR2.

Root MACB mode 0x3 -> 0xd - all other MACB modes 0x3 ( unchanged ).

No timestamps observed from the emulator. Re-set to take timestamps from the VETAR" unit -> all timestamps observed.

10:30

No pulser observed in nn FEE modules 9, 10, 13, 14.

ADC data items check OK - all ~3k - attachment 2

Rates - all OK  - attachment 3

1.8L spectra. - attachment 4,5

1.8W waveforms - attachment 6,7

11:00

Checked daisy chain which goes from 8->14->13->10->9. Replaced the lemo cable from 8->14 and pulsers re-appear. Rates all OK except aida10 which has now disappeared, tried asic check and reset and startup again no change in aida10.

Now majority of pulsers are observable the widths will be noted in a spreadsheet

After replacing lemo for nn pulser chain, all nn pulser widths decreased slightly ~10% for all nn FEEs, little change in pn pulser widths but n pulser for DSSD#1,2 appears more gaussian

11:44

FEE64s power cycled. aida10 now prioducing data.

Removed internal jumper from CAEN N1419ET 

ADC data items check OK - all ~3k - attachment 8

Rates - all OK  - attachment 9

1.8L spectra - attachment 10,11 little change in pn pulser widths but nn pulser widths are slightly broader.

1.8W waveforms - attachment 12,13

15:50 Jumper re-mounted onto caen bias module

ADC data items check OK - all ~3k - attachment 14

Rates - all OK  - attachment 15

1.8L spectra - attachment 16,17

1.8W waveforms - attachment 18,19

16:45 Ground cables attached from water pipe (chamber ground) to lemo cable on aida08 (nn), aida13 (nn), aida09 (nn), aida01 (pn) so that all DSSDs are connected to chamber ground.

ADC data items check OK - all ~3k - attachment 20

Rates - all OK  - attachment 21

1.8L spectra - attachment 22,23

1.8W waveforms - attachment 24,25

 

  482   Tue Jan 30 14:04:30 2024 NH[HowTo] Update MACB Firmware
This is how to update the MACB firmware using the opensource tool xc3sprog


Setup (only has to be done first time)
--------------------
Install xc3sprog and fxload (ubuntu: sudo apt install xc3sprog fxload)

To configure the USB programmer:
Download xilinx.tar.gz and put the hex files in /usr/share

Download the udev rules and put it in /etc/udev/rules.d

Run:
sudo udevadm control --reload-rules

[Taken from https://github.com/timvideos/litex-buildenv/wiki/Xilinx-Platform-Cable-USB-under-Linux]

Plug in programmer and should work

----------------------------------
Programming a MACB

Open MACB side panel and connect JTAG programmer to port 

Plug MACB into NIM bin and power NIM bin on

Run
xc3sprog -c xpc -m -j

Expected output:
XC3SPROG (c) 2004-2011 xc3sprog project $Rev$ OS: Linux
Free software: If you contribute nothing, expect nothing!
Feedback on success/failure/enhancement requests:
        http://sourceforge.net/mail/?group_id=170565 
Check Sourceforge for updates:
        http://sourceforge.net/projects/xc3sprog/develop

JTAG loc.:   0  IDCODE: 0x16d8c093  Desc:                  XC2C128_TQ144 Rev: A  IR length:  8

This confirms the JTAG is working

Then to program  run

xc3sprog -c xpc -m . macb_apr20.jed

The directory should contain the macb_apr20.jed and the xc2c128.map files

Expected output:
XC3SPROG (c) 2004-2011 xc3sprog project $Rev$ OS: Linux
Free software: If you contribute nothing, expect nothing!
Feedback on success/failure/enhancement requests:
        http://sourceforge.net/mail/?group_id=170565 
Check Sourceforge for updates:
        http://sourceforge.net/projects/xc3sprog/develop

Programming row  81                                        
Verify: Success   

Now power off the NIM bin and remove the programmer and put the cover back on
  481   Tue Jan 30 09:14:50 2024 JM, TDTuesday 30 January
09.00 CARME IE514 5.1e-11mbar

09.15 IE514 not working - one of cables disconnected?

      DSSSD #1 bias cable changed from SHV - Lemo 00.250 to SHV - 2x Lemo 00.250

      DSSSD #1 & #2 LK1 removed from n+n Ohmic side bias FEE64s

      aida11 HDMI cabling re-seated at MACB

10.26 DSSSD bias OK - attachment 1
      Ambient temperature +16.2 deg C

10.30 ADC data item stats OK - attachment 2
      aida06 & aida15 > 100k, all others c. 3k consistent with 50Hz pulser

10.34 per FEE64 Rate spectra - attachment 3
       OK except for known issues aida05, aida07 and aida16 - all others OK

      per p+n FEE64s 1.8.L spectra - attachment 4
       pulser peak width 52 ch FWHM aida11- double peaking DSSSD #1 & #2

      per n+n FEE64s 1.8.L spectra - attachment 7
       pulser peak width 77 ch FWHM aida04 - more uniform noise across all DSSSDs

10.42 per FEE64 1.8.W spectra - 20us FSR - attachments 5-6

11:00 LK2 removed from aida01-04. Configuration for all adaptor cards/FEE modules is now the same. No other change to the setup.

12:00 FEEs power on, bias to 150 V. No change in bias observed. 

12.02 ADC data item stats OK 
      aida06 & aida15 > 100k, all others c. 3k consistent with 50Hz pulser

12.07 per FEE64 Rate spectra 
       OK except for known issues aida05, aida07 and aida16 - all others OK
      
12:09  per p+n FEE64s 1.8.L spectra 
       We now observe a large peak around chn 30,000 on all FEEs. Perhaps an effect of some change in the ring?
       pulser peak width 53 ch FWHM aida11- still observe double peaking DSSSD #1 & #2

      per n+n FEE64s 1.8.L spectra 
      Lots of events at ~chn 65,000 - double peaking now observed on aida03 and aida04
      pulser peak width 93 ch FWHM aida14 

12.20 per FEE64 1.8.W spectra - 20us FSR 

      Perhaps bias was off while acquiring and spectra was not cleared properly. Re-do tests

13:40 ADC data item stats OK - attachment 7 
      aida06 & aida15 > 100k, all others c. 3k consistent with 50Hz pulser

       per FEE64 Rate spectra - attachment 8
       OK except for known issues aida05, aida07 and aida16 - all others OK

      per p+n FEE64s 1.8.L spectra - attachment 9 
       pulser peak width 55 ch FWHM aida11- double peaking DSSSD #1 & #2

      per n+n FEE64s 1.8.L spectra - attachment 10
       pulser peak width 65 ch FWHM aida14 - more uniform noise across all DSSSDs

       per FEE64 1.8.W spectra - 20us FSR - attachments 11-12 

      With help from NH updated firmware on all MACB modules to version 4/20 - see https://elog.ph.ed.ac.uk/CARME/482

      Timestamps for all FEE modules now correct - attachment 13

      MAC address of aida07 was found to be 41-b4-16 not 41-d4-16 as was labelled on the FEE module
       re-labelled FEE64
       dhcpd.conf and MIDAS startup.tcl updated, dhcpd and MIDAS 8015 server restarted - aida07 is now working OK
 
      CAEN N1419ET internal jumper fitted - outputs are no longer floating

15:20 ADC data item stats OK 
      aida06 & aida15 > 100k, all others c. 3k consistent with 50Hz pulser - attachment 14

       per FEE64 Rate spectra
       OK except for known issues aida05, and aida16 - all others OK  - attachment 15

      per p+n FEE64s 1.8.L spectra -
       pulser peak width 50 ch FWHM aida11- double peaking DSSSD #1 & #2 - attachment 16

      per n+n FEE64s 1.8.L spectra 
       pulser peak width 67 ch FWHM aida14 - more uniform noise across all DSSSDs

       per FEE64 1.8.W spectra - 20us FSR - aida14 missing from waveforms - attachments 17,18
 
   Swapped the test inputs so that +ve is now from the inverter and -ve direct from pulser.  - attachments 19,20. Observed reduced noise for when plugged directly into the pulser than when connected from the inverter. 



   Lowered threshold from 0x64 (1 MeV) to 0x20 (320 keV) - attachment 21, 22 - Observed increase in rates, in particular for the DSSD #1,2 pn side. Conclusion -> increased pulser width for DSSD #1,2 is real and not from the pulser
 


   aida01 lemo to water pipe (chamber ground) connection installed 

   ADC data item stats OK 
      aida06 & aida15 > 100k, all others c. 3k consistent with 50Hz pulser 

       per FEE64 Rate spectra
       OK except for known issues aida16 - all others OK 

      per p+n FEE64s 1.8.L spectra -
       pulser peak width 48 ch FWHM aida11- double peaking DSSSD #1 & #2 - attachment 23

      per n+n FEE64s 1.8.L spectra 
       pulser peak width 67 ch FWHM aida14 - attachment 24

       per FEE64 1.8.W spectra - 20us FSR - slightly reduced 100 kHz on pn waveforms - attachments 25,26

   Disconnected aida06 and replaced with the spare FEE module (mac address: 42:0d:16). Power cycled and updated config files. 

   ADC data item stats OK 
      aida15 > 100k, all others including aida06 c. 3k consistent with 50Hz pulser - attachment 27

       per FEE64 Rate spectra
       OK except for known issues aida16 - all others OK - attachment 28

      per p+n FEE64s 1.8.L spectra -
       pulser peak width 48 ch FWHM aida11- no longer observe double peaking DSSSD #1 & #2 but the pulser peaks are still broader compared to DSSD #3,4 - attachment 29

      per n+n FEE64s 1.8.L spectra 
       pulser peak width 55 ch FWHM aida14 - attachment 30

       per FEE64 1.8.W spectra - 20us FSR - greatly reduced 100 kHz on pn waveforms, dodgy FEE previously aida06 must@ve had some impact on DSSD#1,2 - attachments 31,32
      
  480   Mon Jan 29 15:32:08 2024 TD, JMMonday 29 January contd.
Test + and test - daisy chain cabling extended from DSSSD #1 to #2 to DSSSD #1 to DSSSD #4
 terminated by 50 Ohm at end of daisy chain

Grounding cabling extended from DSSSD #1 ( only ) to DSSSD #1 to DSSSD #4
 grounding cables daisy-chained between 4x adaptor PCBs of each DSSSD - no inter-DSSSD grounding cables

DSSSD#1 bias cable SHV-Lemo 00.250 ( to p+n bias adaptor PCB )
DSSSD#2 - DSSSD #4 bias cables SHV - 2x Lemo 00.250 ( to p+n/n+n bias adaptor PCBs )

LK1 fitted for DSSSD #1 & #2 n+n Ohmic side bias adaptor PCBs
LK2 not fitted on any adaptor PCBs? to be checked

CAEN N1419ET internal jumper fitted ( non-floating outputs )? to be checked

16.54 DSSSD bias OK - see attachment 1

      FEE64 temperatures - attachment 2
       aida06 ASIC temp low

      System wide checks and WR - attachments 3-6
       global clock status aid04 0x6, aida06 0x4
       multiple FPGA timestamp and WR decoder errors for MACBs 4 and 5 ( of 5 )

       WR timestamps OK for aida01-aida08 ( w/ MACB firmware version 04/20 )

     ADC data item stats - attachment 7
      all rates consistent with 50Hz pulser *except* 
       aida11 global clock fail
       aida07 no boot
       aida16 no power
       aida06 ? 
       aida15 faulty asic #2

     per FEE64 Rate spectra - attachment 8

     per p+n and n+n FEE64 1.8.L spectra - attachments 9-12
      Pulser peak width p+n side 52 ch FWHM ( aida12 ), n+n side 70 ch FWHM ( aida14 )

     per p+n and n+n 1.8.W spectra - 20us FSR - attachments 13-16
  479   Mon Jan 29 08:52:07 2024 TD JMMonday 29 January
09.50 CARME IE514 5.1e-11 mbar ( ambient temperature 16 deg C )

10.24 DSSD bias OK - attachment 1

10.40 MACB configuration

      1              2      3      4      5

mode  0x3            0x3    0x3    0x3    0x3

      from VETAR2    from 1 from 1 from 1 from 1
      to 2           aida01 aida05 aida09 aida13
      to 3           :      :      :      :
      to 4           :      :      :      :
      to 5           aida04 aida08 aida12 aida16

WR emulator cables connected to MACB #1 rear panel

DAQ reset and DAQ setup between each test below - the FEE64s were NOT power cycled between tests


      Per https://elog.ph.ed.ac.uk/CARME/478 test WR

      connecting/disconnecting HDMI cable from VETAR2 to MACB level 0 input was tested 20.1.24 per https://elog.ph.ed.ac.uk/CARME/477

      With VETAR2 abd all MACB mode sttings 0x3 observe same WR issues as previously

      Noted clock & sync cables from WR emulator were swapped
      ( pay attention to SMA cables cf. Zybo board *not* NIM module as shown in https://elog.ph.ed.ac.uk/DESPEC/517 )

      Change MACB level 0 mode from 0x3 to 0xd - all other MACBs mode 0x3
      HDMI cabkle from VETAR2 to MACB level 0 input disconnected

      System wide checks - global clocks all fail - see attachments 2-5


11.14 Reduce MAC config to MACB level 0 plus one MACB connected to FEE64s aida01-aida4
      All MACB modes 0x3

      System wide checks aida01-aida04 OK. Same problem observed with WR timestamp for aida01-aida04

      See attachments 6-9

11.22 Reduce MAC config to one MACB connected to FEE64s aida01-aida4
      MACB mode 0x3

      System wide checks aida01-aida04 OK. Same problem observed with WR timestamp for aida01-aida04

      See attachments 10-13

11.26 One MACB config using MACB 2 of 5 in NIM bin connected to FEE64s aida01-aida4
      MACB mode 0x3

      System wide checks aida01-aida04 OK. Same problem observed with WR timestamp for aida01-aida04

      See attachments 14-17

11.36 Connect VETAR2 output directly to aida09

      System wide checks for aida09 OK.Valid WR timestamp appears to be generated.

      See attachments 18-20

      Check NIM bin with MACBs etc
        +/-6V, 12V and 24V lines OK

11.42 Connect VETAR2 to MACB 1 of 5 ( mode 0x3 )
      Connect aida09 to MACB 1 of 5
      Global clock OK. WR decoder status error 0xd reported for aida09. Valid WR timestamp for aida09.

      See attachments 21-24

11.53 Connect VETAR2 to MACB 1 of 5 ( mode 0x3 )
      Connect aida09-aida12 to MACB 1 of 5
      Global clock OK. WR decoder status error 0xd reported for aida09-aida12. WR timestamps for aida09-aida12 very different - aida10 might be correct timestamp?

      See attachments 25-28

11.53 Connect VETAR2 to MACB 2 of 5 ( mode 0x3 )
      Connect aida09-aida12 to MACB 2 of 5
      Global clock OK. WR decoder status error 0xd reported for aida09-aida12. WR timestamps for aida09-aida12 very different - aida10 & 11 approx correct but 11 has 
lower timestamp than 10 so probably not quite right

      See attachments 29-32


15.00 Per https://elog.ph.ed.ac.uk/CARME/478 systematically test 1x MACB configurations to find 'working' MACBs

      Identify 3x MACB modules which produce sensible WR timestamps - all labelled ' firmware version 04/20 '.
      The other 5x MACBs are either unlabelled or labelled unknown.

      MACB configuration now uses 04/20 MACBs for MACBs 1-3 and unlabelled 2xMACBs fpr MACBs 4-5
      aida01-aida08 should produce WR timestamps OK

      Will attempt to update firmwre of 5x MACBs tomorrow pm with NH
  478   Sat Jan 27 20:38:15 2024 PJCS[INFO] Steps to diagnose White Rabbit problems

This is a suggested sequence of actions to help with the current problem with the White Rabbit timestamp distribution. Test the system between each intervention. If a test provides an answer to the problem then proceed accordingly rather than continue with the suggestions.

  • Please sketch the current MACB wiring.
  • Unplug and replug the connectors of the VETAR HDMI connecting cable. If possible clean the pcb and cable HDMI connectors with contact cleaner.
  • Disconnect the VETAR and connect the Emulator. If the clocks are not correct then at the Local and Waveform control page change the entry for LMK03200 control register at offset 5 to 0xB. If clocks are now ok then repeat for all the FEEs. Be aware that this value will be set back to 0x1B during SETUP. If the Emulator fails then continue with VETAR and setting 0x1B.
  • Connect one or more FEE64 to the base MACB.
  • Change the base MACB and repeat the test above.
  • Connect a FEE to the VETAR direct. 

I am at present unable to advise further. However should there be more then I will update this entry.

  477   Mon Jan 22 16:29:31 2024 TDSaturday 20 January
11.00 No WLAN/WiFi
      came-gsi private LAN appears to be OK

      AIDA FEE64 power ON

11.09 FEE64 temps - attachment 1
      aida06 low, otherwise OK

11:10 CARME IE514 4.0e-11 mbar, temp unknown
      IE514 OFF

11.22 Detector bias ON - attachment 2
      found ch 2 up ramp set to 63V/s! changed to 1V/s

11.30 system wide checks OK except 
      aida02 & aida06 fail WR decoder status

      WR ts control - attachments 3-4


12.02 ASIC settings updated for aida09-aida16
      ASIC settings DB key 2024Jan20-12:02:08

      Unable to save DB key to Options - use ASIC settings restore as workaround

12.30 Layouts for 16x FEE64s created - directory ~/LayOut/16xFEE64
      Note layouts include spectra from aida07, aida08 and aida16 so an error message will be generated but it can be ignored


12.40 per FEE64 Rate spectra - attachments 5-6
      per FEE64 p+n & nn 1.8.W spectra - attachments 7-8

12.57 per FEE64 p+n & n+n 1.8.W spectra 20us FSR - attachments 9-10

Test signals 

EG&G Ortec 448 Pulser
Amplitude 90,000
attenuation x10
Normalise 5.00
t_r 100ns
tau_d 50us
Frequency 50Hz ( relay line freq )
Polarity +

test + direct from 448, test - from Cooknell SA1 Sum & Invert 

Test + daisy chain -> aida01 - aida02 - aida05 - aida06 - 50 Ohm
Test - daisy chain -> aida03 - aida04 - aida07 - aida08 - 50 Ohm

aida08 ( DSSSD #2 n+n Ohmic bias ) LK2 removed, LK1 added
aida04 ( DSSSD #1 n+n Ohmic bias ) no LK2 fitted, LK1 added

DSSSD#1 SHV - 2x Lemo 00.250 cable replaced by standard SHV - BNC/BNC Lemo-00.250 coaxial cable

Adaptor PCB ground daisy chain aida01 - aida02 - aida03 - aida04

DSSSD#2 - DSSSD#4 HV cabling unchanged

15.30 Power ON
      System wide checks OK except
          aida02 fail WR decoder status - attachment 11

      FEE64 temperatures - attachment 12
       aida06 ASIC temp low

      DSSSD bias - attachment 13
     
      ADC data item stats - attachment 14
       slow comparator 0x64

      per FEE64 Rate and 1.8.W 20us FSR spectra - attachments 15-17

16.00 per FEE64 rate spectra - attachment 18

16.15 per FEE64 Rate and 1.8.W 20us FSR spectra - attachments 19-21

17.36 Switch from VETAR2 ( all MACB modules mode 0x3 ) to NIM WR emulator ( root MACB module 0xd, all others 0x3 )
      All FEE64s fail Global clock and WR decoder - attachments 22-24
      WR timetsamp control - attachment 25


18.10 Noted MACB #4 mode actually set to 0x2 *not* 0x3 - changed to 0x3

      Re-test with VETAR2 WR - same behaviour as observed this morning - see attachments 26-30




Summary

n+n sides appear to be OK but noisy

no data observed from aida06 & aida11
 aida11 known to be disconnected from adaptor PCB

aida15 asic #2 may be faulty

aida01, aida05, aida12 - data suggest FEE64/adaptor PCB alignment issue ?

aida02, aida03, aida09, aida14 - data suggest adaptor PCB damaged pins ?


aida07 - network config issue ?

aida08 - not completing boot

need to attach sys console to aida07 & aida08 to diagnose problem.
  476   Fri Jan 19 11:17:50 2024 JG, TDFriday 19 January
12.20 CARME IE514 4.1e-11mbar, ambient temperature 16.7 deg C

      FEE64 cabling complete

      WR source - VETAR2 borrowed from Nik Kurz and opwerating standalone ( no BuTIS reference )
      All MACBs mode 0x3


15.20 Power ON FEE64s #1 - #15

      aida02 did not boot - /etc/dhcp/dhcpd.conf typo - fixed

      aida04 did not boot - AIDA FEE64 PSU output #4 fuses open circuit - fuses replaced from AIDA FEE64 PSU spare channel ( aida16 not connected yet ) - fixed

      aida07 did not boot - power OK - FEE64 comfort LEDs OK - network switch link LED OK - network config?

      aida08 boots - power OK - comfort LEDs OK - network switch link LED OK - requests RFS / but does not request RFS MIDAS - can ping aida08 OK, cannot telnet aida08 'connection refused'

      aida09 did not boot - FEE64 comfort LEDS OFF - swap FEE64 #9 & #10 power cables - aida09 boots OK, aida10 does not boot
      Restore #9 & #10 power cables - swap AIDA FEE64 outputs #1 & #2 ( aida09 & aida10 ) - aida09 boots OK
      Conclude AIDA FEE64 PSU ( no label ) output #1 is faulty. Replace AIDA FEE64 PSU ( label ISOL3 ). aida09 & aida10 boot OK

      aida15/aida16 MAC addresses were swapped in https://elog.ph.ed.ac.uk/CARME/467 - fixed

      aida01-aida06, aida09-aida15 boot OK - aida07 does not boot, aida08 does not complete boot

16.25 Configure MIDAS DEE64 servers aida01 - aida06, aida10 - aida15 

      FEE64 temperatures - attachment 1
      aida06 looks low ( c. 25 deg C cf. c. 40 deg C )

      System wide checks - attachments 2-8
       aida02 fails WR decoder

      WR timestamps ( attachments 7 & 8 ) look odd/low - values always c. 0x2???


17.30 CARME IE514 OFF

      WR NIM emulator SMA cables to root MACB ( rear panel ) disconnected at front panel

      Power cycle FEE64s

      System wide checks & WR timestamp control - attachments 9-14
       no change

17.45 Detector bias -20, -100 & 150V - attachments 15-17
      CAEN N1419ET non-floating outputs ( to be checked ), HV cable SHV - 2x Lemo 00.250, LK1 not fitted 

18.40 Detector bias -150V attachment 18




To Do
Check ASIC settings
Configure HV
Check LKs
Ground cabling
      
  475   Thu Jan 18 17:21:00 2024 TDThursday 18 January
Network ( dhcpd.conf ) and MIDAS startup.tcl/Options/ASIC settings/rfs filesystems configured


17.20 CARME IE514 4.1e-11mbar, ambient temperature 16.7 deg C
      Pressure appears to be drifting upwards - see attachment 1
  474   Wed Jan 17 19:21:20 2024 TDAIDA spares
Spare AIDA hardware 

1x AIDA FEE64 PSU
1x USB-controlled ac mains relay 

Transferred from EX 2.039 store ( ESR roof ) to CARME storage bins adjacent to CRYRING counting room ( CRYRING roof ).

CRYRING Hall

12 USB-Jtag cables ( sys console ) + 1x to Edinburgh Jan 2024 to fabricate more with spare Jtag PCBs
7x HDMI FEE64 - MACB cables
0x 0.3m HDMI-HDMI cables - require 8x 0.3m HDMI - HDMI cables + 2x 3m HDMI-HDMI cables
2x MACB
1x NIM WR emulator
1x Dlink 7 port USB extender
1x Netgear 24-port Gbit network switch
1x CAEN N1419ET ( ch #0 u/s ? )
0x 10m RJ45 Cat 6 network cables yellow/green/purple/brown - require 2x of each
  473   Wed Jan 17 19:15:47 2024 TDWednesday 17 January
CAEN N1419ET remote control via carme-gsi and nnrpi1 - see https://elog.ph.ed.ac.uk/CARME/472

15x ( of 16x ) AIDA power cables available, labelled #1 - #15, installed

16x ( of 16x ) RJ45 Cat 6 network cables available, 15x labelled #1 - #15

16x ( of 16x ) HDMI cables available, labelled #1 - #16
  472   Wed Jan 17 10:31:05 2024 TDRemote control of CAEN N1419ET from nnrpi1
Login to nnrpi1 from carme-gsi using ssh

Execute the caen_screen.sh script

pi@nnrpi1:~ $ cat ./caen_screen.sh
if ! screen -list | grep -q "caenhv"; then
	screen -f -S caenhv /dev/ttyACM0 9600
else
	screen -x caenhv
fi
pi@nnrpi1:~ $ ./caen_screen.sh

Enter the text string 'caen'

If the CAEN display program is already running you will only see page updates/changes.
- enter q ( quit ) and then d ( display ) to re-display the entire screen.

See attachment 1
  471   Tue Jan 16 16:01:48 2024 TDChange CARMEraspberrypi to nnrpi1
Change RPi from CARMEraspberrypi to nnrpi1 to provide carme-gsi access to Rly16 to control USB-controlled ac mains relay.
nnrpi1 is already defined in carme-gsi:/etc/dhcp/dhcpd.conf

Transfer RPi network connection from CRYRING local network to AIDA 24-port network Gbit switch

Change /etc/hostname from

CARMEraspberrypi

to

nnrpi1

Change /etc/hosts from

127.0.0.1	localhost
::1		localhost ip6-localhost ip6-loopback
ff02::1		ip6-allnodes
ff02::2		ip6-allrouters

127.0.1.1	raspberrypi	

to

127.0.0.1	localhost
::1		localhost ip6-localhost ip6-loopback
ff02::1		ip6-allnodes
ff02::2		ip6-allrouters

127.0.1.1	nnrpi1

Reboot RPi

Login to RPi and start MIDAS 8015 service

/MIDAS/TclHttpd/linux-arm/TclHttpd-server &

Open new browser tab on carme-gsi

http://nnrpi1:8015/AIDA/Rly16/Rly16.tml
  470   Tue Jan 16 14:36:30 2024 TDCARME configuration
FEE64

#1 - #16 -> aida01 - aida16

Network configuration - /etc/dhcpd/dhcpd.conf - attachment 1

MIDAS Options

Added Options sub-directories for aida09-aida16 - attachment 2

MIDAS ASIC settings

Added sub-directories for aida09-aida16 - attachment 3

MIDAS configuration - /MIDAS/config/TclHttpd/carme-gsi"8015 - attachment 4
Currently only aida01-aida08 defined

MIDAS embedded rfs filesystems for aida01-aida24 alreday exist - attachment 5

All FEE64s connected to rev B CARME adaptor PCBs
LK1 fitted to n+n bias FEE64 ( aidaxx, aidaxx, aidaxx, aidaxx )
No other LKs fitted


MACB

WR source NIM WR emulator or VETAR2

See https://elog.ph.ed.ac.uk/DESPEC/517 for MACB mode configuration


CAEN N1419ET


EG&G Ortec 448 Pulser


ASIC settings


Options
  469   Tue Jan 16 08:57:44 2024 JG, DR, TDTuesday 16 January
10.00 Outflow manifold returned from workshop with leaking attachment point welding re-worked.

11.30 Outflow hoses re-attached to outflow manifold. Faulty male Swagelok replaced.
      Outflow manifold tested on floor - no leaks observed.

      Mounted outflow manifold. All inflow/outflow hoses attached to FEE64s - see attachments 1-4.
      FEE64 #11 detached adaptor PCB - temporarily secured by cable tie.

      Inflow/outflow valves opened. Pin hole leak observed at 3/8" connector gasket - indicated by arrow attachment 
3.
      Leak fixed by tightening fitting.

      1x spare inflow/outflow connected together - attachment 4.

11.40 CARME IE514 3.8e-11mbar, ambient temperature 16.5 deg C

15.05 8x AIDA FEE64 power cables ( #1 - #8 ) installed FEE64s #1 - #8
      Note - Samtec lock clips broken on 3 of 8 cables due to careless handling/poor storage
      AIDA FEE64 power cables tied to one of the FEE64 copper coolant pipes.

      See attachments 5-7

15.30 USB-controlled ac mains relay - USB connected to RPi CARMEbakingpi currently used for CAEN N1419ET 
      Rear panel key switched to override - no flow/d.p. sensors currently installed - front panel water ( yellow ) and override ( red ) LEDs active.

16.15 Transfer CARMEbakingpi from CRYRING network to AIDA 24-port Gbit network switch connected to carme-gsi
      CARMEbakingpi becomes nnrpi1 p- can connect OK via ssh from carme-gsi

      Start server ( port 8015 ) manually

      ssh pi@nnrpi1
      
      /MIDAS/Tcl/Httpd/linux-arm/Tcl-Httpd-server &

      See attachment 8
  468   Mon Jan 15 18:51:28 2024 JG, DR, TDMonday 15 January
Cooling water manifolds


The manifolds were returned this afternoon.

Inflow manifold - blue hose, female Swagelok connectors

Attached to inflow hose and bled using bleed/drain valve.
Small leak observed at the 3/8" connection between the manifold and the bleed/drain valve - attachment tightened and leak fixed.
Inflow manifold installed.

Outflow manifold - red hose, male Swagelok connectors

Attached to outflow hose and bled using bleed/drain valve.
Small leak observed at the weld immediately *below* the weld that had been re-worked - this weld had not been re-worked.
Significant leak observed at one of the male Swagelok connectors in the lower half of the manifold - it does not seal correctly.


DR will return this manifold to the workshop for welding re-work Tuesday am.
The faulty male Swagelok will be replaced.


AIDA cabling


16x AIDA PSU, HDMI and RJ45 Cat 6 cables and 4x SHV-BNC cables required.

Available

8x AIDA PSU cables 
1x AIDA PSU cable ( in one of two white plastic storage boxs with 7x FEE64s returned from STFC DL ) now appears to be missing

16x AIDA HDMI cables 

8x RJ45 Cat 6 network cables
1x RJ45 Cat 6 network cable ( network switch to CRYRING network patch panel - to be replaced )
3x RJ45 Cat 6 network cables found in AIDA plastic box ( S4 Messehutte )

3x SHV-BNC 

Will check DESPEC store above ESR with NH tomorrow afternoon.
  467   Sun Jan 14 12:31:32 2024 JM, CBFEE positions mac addresses

*Updated following FEE changes. Correct as of 31/01/24

FEE modules are labelled according to the diagram in attachment. Note attachment is looking downstream. Each FEE module is also labelled with its assigned number and whether it is a bias connection on the FEE holders. MAC addresses and FEE descriptions are as follows;

1 -  41:d8:20  - top left detector -  pn bias

2 -  41:d8:2a  - top left detector - 

3 -  41:d0:05  - top left detector - 

4 -  41:b4:0b  - top left detector -  nn bias

5 -  42:0d:0c  - top right detector -  pn bias

6 -  41:0d:16  - top rightdetector - 

7 -  41:b4:16  - top right detector - 

8 -  41:ba:2b  - top right detector -  nn bias

9 -  41:d0:0e  - bottom left detector -  nn bias

10 -  41:d8:2b  - bottom left detector - 

11 -  41:d7:cc  - bottom left detector - 

12 -  41:d7:c2  - bottom left detector -  pn bias

13 -  41:ba:22  - bottom right detector -  nn bias

14 -  41:ee:72  - bottom right detector - 

15 -  41:a0:71  - bottom right detector -

16 -  41:ee:71  - bottom right detector -  pn bias

  466   Sat Jan 13 14:32:21 2024 CB, JMLeaky manifold and cable management

Found two leaks in the welding of the manifolds, marked them in tape. Manifold to be repaired on Monday. Remember to re-install missing hoses, which are  in a bucket in the cave.

See attached how it is supposed to look like.

Left hand hose (inlet) was too short and was cut to dismount the manifold. Davide will obtain a new tube of the same type and mount it for us. NOTE THAT THE INLET connector (BLUE manifold) was leaking. Unclear if this was the hose or poor plumbing.

Installed cable management system that was never used during the last beamtime. Started to thread cables through it. There should be sufficient space to re-install the manifolds with the cable tray in place. If this is difficult, the tray is installed with zip ties only so altering position should be relativly straightforward. The mounting points of the outlet manifold on the CARME chamber can also be rotated around to give a bit more space.

Cleaned up some boxes and moved heaters and some empty boxes to the CRYRING roof.

ELOG V3.1.4-unknown